Kyla wants to save $8000 in 10 years she has three options determine the total amount that Kyla most deposit for each option
monthly deposits paying 3% interest compounded monthly
sn=a(r^n-1)/r-1
8000(1.0025^120-1/1.0025-1
8000x(0.349353/0.0025
8000x139.74
8000/139.74=x
x=57.25
quarterly deposits paying 3% interest compounded quarterly
sn=a(r^n-1/r-1
8000(1.0075^30-1/1.0075-1
8000x0.2512717/0.0075
8000x33.50
8000/33.50=x
x=238.80
I believe that the problem is in order. Though, i feel that my answer is coming up short. Can someone help explain what im doing wrong?

Answer by ankor@dixie-net.com(15746)

: Kyla wants to save $8000 in 10 years she has three options determine the total amount that Kyla most deposit for each option
monthly deposits paying 3% interest compounded monthly
sn=a(r^n-1)/r-1
8000(1.0025^120-1/1.0025-1
8000x(0.349353/0.0025
8000x139.74
8000/139.74=x
x=57.25
I check this using the annuity feature on the Ti83, and it is correct
:
quarterly deposits paying 3% interest compounded quarterly
sn=a(r^n-1/r-1
8000(1.0075^30-1/1.0075-1
8000x0.2512717/0.0075
8000x33.50
8000/33.50=x
x=238.80
On this one I think you made a mistake, it should be 40 periods, you have 30
x=172.24